OpenStudy (anonymous):

Multiply and simplify: ((x^(2)-49)/x) * ((x+6)/(x+7))

OpenStudy (anonymous):

\[\frac{ x ^{2}-49 }{ x } * \frac{ x+6 }{ x+7 }\]

OpenStudy (anonymous):

what do the x stand for ?

OpenStudy (anonymous):

I factored it out to be \[\frac{ (x+7) (x-7) (x+6) }{x(x+7) }\]

OpenStudy (anonymous):

And then divided out the (x+7)'s s that it's \[\frac{ (x-7) (x+6)}{ x }\]

OpenStudy (anonymous):

but is this the simplified form? can i simplify it further?

OpenStudy (anonymous):

i know but what does the x stand for

OpenStudy (anonymous):

x is just a variable. I don't have to figure out what x is. I just have to simplify the equation

OpenStudy (anonymous):

@phi Sorry to bother you, I just want some help!

OpenStudy (phi):

that is as simple as it gets

OpenStudy (anonymous):

Thank you.